Crew 215 Science Report 16Nov2019

[title Science Report – November 16th]

[category science-report]

Science Report 16 November 2019
SOL: 6
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The remaining micrometeorite samples collected from yesterday
were processed and packaged in the morning. In addition, gypsum was ground for cement
experimentation with aggregate samples (fine sand, coarse sand and pebbly sand) the
only outstanding components. In the afternoon, more micrometeorite samples were
collected and transferred to the science dome.
Green Hab: Green Hab operations have continued with a number of herb and vegetable
species now sprouting. See Green Hab report.
RAM Operations: N/A
EVA: The sixth EVA for science saw us returning to the MM5 (518462E 4248949N) site
where the remaining micrometeorite samples were collected. 100% of the grid has now
been sampled and the grid retrieved. See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin(TM)
monitoring vests. The simulated emergency retrieval of the engineer has provided a
great deal of data for formulating procedures to assist those in remote and isolated locations.
Additional activities: N/A.

Copyright © The Mars Society. All rights reserved. | Main Site